5. Will this profit function have a maximum, if so, what is it?
To find the maximum, I will use the formula;
P(x) = -.1x2 + 50x - 300, where x is the number of items sold, x ≥ 0.

these are the numbers to use in this function
90 34419
80 33936
70 33451
60 32964
50 54975

Answer by stanbon(75887)

Will this profit function have a maximum, if so, what is it?
To find the maximum, I will use the formula;
P(x) = -.1x2 + 50x - 300, where x is the number of items sold, x ≥ 0.
----------------
Since the coefficient of x^2 is negative P(x) has a maximum value
at the vertex.
----
Vertex occurs where x = -b/(2a) = -50/(2*-0.1) = 250
----
Max value = P(250) = 5950
=============
Cheers,
Stan H.